INFRARED EMISSION SPECTRA OF MOLECULAR IONS IN FAST ION BEAMS

\underline{M.D. TISSANDIER}, A.D. EARHART, and J.V. COE, Department of Chemistry, The Ohio State University, Columbus, OH 43210-1173..

Detection and dispersion of infrared emission spectra from collisionally activated fast ion beams has been previously demonstrated. A new apparatus has been constructed based on an electrostatic particle guide. Progress regarding adaptation of the device to a commercial FTIR and development into a general method for recording low resolution infrared spectra of molecular ions will be discussed.